Prev Next

SW&IT Meeting Minutes 2021-11-01

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211018
LR20211018 - Bec, Matthieu 80.0
LR20211018 - Mechtley, Brandon 80.0
LR20211018 - Eychaner, Glenn 40.0
LR20211018 - Edmeades, David 40.0
LR20211018 - Grenz, Paul 12.0
LR20211018 - Hooper, Stephen 40.0
LR20211018 - Zhang, Yang 80.0
LR20211018 - Zhang, Xianyu 93.0

News from management

  • AO group re-arrangement (LF departure): GB move to ENG, XZ move to SWG, DM, JC, GT report to CV
  • DC part-time for Steward
  • UoA vaccination requirements
  • no news from board about new Director and search
  • return your PA (thanks to those who did already)

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
    • ATT/SW one-slide recap for the board
    • close swg recruitment
    • msgd/rtdb fedora/python3 for DMS, WIP
  • This week
    • Performance Appraisals
    • git/github "policy" (= philosophy slides) for new staff

YZ

  • Previous:
    • LBC alert list and MODS WCS program discussion with Olga.
    • RM507 machine setup meeting on Monday.
    • LBC controller swap (#4 -> #1, then back to #4), bias level measurement and adjustment (with Olga and Alessandro).
    • Argos swingarm testing/debugging with Paul.
    • TMS software share (with GMT) discussion??
    • Information Security Awareness Certification.
  • Next:
    • Finish LBC alert list.
    • Continue MODS WCS work with Olga.
    • TMS meeting and software work.
    • Argos swingarm testing/debugging with Paul.

DE

  • Windows server ping. Again. I'm trying to impress a sense of urgency, but communication is nonexistent from them. Escalation?
  • Util2 threw a wobbler on Friday and is very unhappy. I will open a ticket with Dell. I'll probably have to go in and poke at it to troubleshoot.
  • Related--the util2 failure revealed some missed DNS server configs; most if not all of those are now resolved.
  • Tucson Synology SSD oops. I couldn't find an exact model match so I got a different brand. Turns out SSDs are still variable in actual capacity despite being labeled "X TB". I thought we were past that but apparently not. Cindy will return it and I'll get a new one.
  • New Synology buy
  • Obs1 went odd, restarted Xwindows which solved most of the issue but ended up needing a reboot to get audio back.
  • Rm507-1 had a video failure with all screens flashing on and off. Hard reboot fixed. Probably video driver issue?
  • Monitor.mountain DB error, rebooted.

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Discussion with Dave Thompson regarding the unstable amplifier 2 in LUCI2 (IT 6818). It appears we may need a more sophisticated solution than the current fixed value multiplier.
    • More discussion of LUCI FITS headers, stale DD values, and AO.
      • The AO team was informed that they need to be responsible for keeping DD values up to date.
    • Fixed the LUCI simulation configuration on osa-luci.
    • Finally, after more effort than expected, ran Aladin 11 on all the simulators. Unfortunately, it does not behave appreciably better than Aladin 10.
      • I also ran Aladin 10 and 11 separately from the RTD, and did not find any of the problems when running commands manually.
      • I now have pages and pages of exception stack traces to look at in order to fix the RTD/Aladin problems. One issue may be that the RMI server listeners call Aladin directly with commands.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Evaluate whether we should consider moving to DS9. After running DS9 in my simulation, it has much better performance than Aladin, but may not have all the features we need.
    • Continue working on the RTD issues.

SH

  • Vacation
  • DNS bounce, DNS things

next:

  • IT spending
  • obs computer stuff from meeting
  • monitor.mountain issues
  • Mark computer stuff tomorrow, Steward visit tomorrow.

BM

  • Last week:
    • LBTI engineering night. Had some issues with the install, auxloops not installed. Conf/calib missing files
    • Proposed code deployment changes to AO group. Will go for it this week.
    • Merged latest changes (gain ramp, Python formatting, GCC warning tweaks), still need to be tested.
    • Met with Gustavo, Alessandro. Got the details on the OCAM lab test setup. Might see if EngPanel can be run from virtualbox on spare-soul-wfs so we don't need the laptop.
    • Looked over HouseKeeper, DiagnApps.
  • In the queue:
    • HouseKeeper logging during SwitchBCU timeout. Know what I need to do, just need to implement it.
    • Carefully make ao-supervisor installs consistent.
    • AO data for LUCI.

XZ

  • Last week
    • DX ASM data analysis with last week test in the clean room with Guido
    • DX ASM data analysis with past shell rested telemetry with Guido
    • DX ASM data documentation with Guido
    • OCAM lab test setup, AO SW/plan, wind speed/TSS issues, etc
  • This week
    • DX ASM path trade-off study with Guido
    • Fault recovery dicusstion/implemention with Guido and Arcetri
    • AO support

PSG

  • Last week:
    • Testing swing arms for ARGOS. No issues, The problem we saw last time did not re-occur.
    • Worked through GUI display of status. Display issues are fixed now as well.
  • This week:
    • Make change to HL2000 permananent (more than 1 connections).
    • Test to ensure it works.
Topic revision: r13 - 01 Nov 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback